What is the solution set of the inequality below?

17-|x+6|<-17


Answer by dabanfield(803)

17-|x+6|<-17
From above we have:
|x-6| > 34
When (x-6) >= 0 (i.e. x >= 6)we have:
|x-6| = +(x-6) = x-6 so:
We have then when x >= 6, x-6 > 34
The intersection of these two solution sets is:
x >= 6 and x > 40 or
x > 40.

On the other hand:
When (x-6) < 0 we have:
|x-6| = -(x-6) = -x + 6
So when x-6 < 0 we have x < 6 and
-x + 6 > 34 or
x < -28
The intersection for these two solution sets is:
x < 6 and x < -28 which is x < -28.
Putting the two solution sets together we have x < -28 or x > 40.